OpenStudy (anonymous):

The ratio of e/m of aproton to that of an alpha particle is 1:1 , 1:4, 1:2, 1:1/4

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

An \(\alpha\) particle is 2 protons and 2 neutrons.. not sure what "e/m" means

OpenStudy (anonymous):

Caherge to mass ration is e/m .

OpenStudy (anonymous):

Charge *

OpenStudy (aaronq):

i've never seen "e" as charge. it's usually "z", or "q".

OpenStudy (anonymous):

@zpupster Solve it please ..

OpenStudy (aaronq):

you don't need to use math, it's simple. Mass of proton 1, mass of alpha particle 4 same charge for both

OpenStudy (anonymous):

How both have Same charge ?

OpenStudy (aaronq):

lol i mean, the \(\alpha\) particle obviously has a 2+ charge

OpenStudy (zpupster):

Worked Example What is the charge to mass ratio of a helium nucleus (alpha particle)? We will assume that the proton and the neutrons have the same mass, 1.67 × 10-27 kg. In a helium nucleus, there are 2 protons and 2 neutrons, making up 4 nucleons. Charge to mass ratio = 2 × 1.6 × 10-19 C 4 × 1.67 × 10-27 kg = 4.85 × 107 C kg-1 http://www.antonine-education.com/Pages/Physics_1/Particles/PP01/Particles_1.htm

OpenStudy (aaronq):

\(\dfrac{\dfrac{1}{1}}{\dfrac{2}{4}}=\dfrac{1}{\dfrac{1}{2}}=2:1\) ? it makes sense because the proton is 1:1, and the alpha p is 2:4, so it's half, though it's weird that that answer is not given.

OpenStudy (zpupster):

yep 1:2

OpenStudy (anonymous):

Ty :D